Carly is 5 years older than Dan.In 10 years the sum of their age will be 75.Find the age of each man now

Answer by ikleyn(52803)

.
C - D = 5.             (1)
(C+10) + (D+10) = 75.  (2)


It is equivalent to

c - D =  5,            (3)
c + D = 55.            (4)

Now add the equations (3) and (4) and see what will happen.

Answer. C = 30, D = 25.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".